Grade 12th passAlgebra

What is the remainder obtained by dividing 2^1024 by 5

Michu
8 Years agoGrade 12th pass

1 Answer

Nikhil
ApprovedApproved Tutor Answer8 Years ago
dear friend,
as we go following sequence
2/5  rem. =2        4/5  rem.=4           8/5 rem.=3          16/5 rem.=1
32/5 rem. =2    
so you can see that the sequence of reminders star repeating after 4 powers 
so simply divide 1024 by 4    reminder comes out 0
so the reminder which comes out on dividing 2^0 = 1 by 5   will be same that in case of 2^1024 dividing by 5 i.e.  1
so correct answer is 1